# FakET: Simulating Cryo-Electron Tomograms with Neural Style Transfer

### v1

@@ 3 Methods
The methods described in this section represent surrogate models that mimic the behaviour of the aforementioned SHREC physics-based simulator. The methods are to be applied in the projection space, i.e. all of them take noiseless (and ice-less) projections as inputs, shown in Figure 2, and produce real-like looking projections by matching their appearance with a target “style” projections. The outputs are ready to be reconstructed into final tomogram volumes using a reconstruction algorithm of choice.
There are multiple reasons why we decided, and why it is reasonable, to model the simulator’s behaviour in projection space. Firstly, different reconstruction algorithms produce different types of artifacts and are suitable in different situations. Therefore, simulating only projections means the models do not have to mimic the artifacts crated by the reconstruction algorithm (e.g. smearing due to the missing wedge). Moreover, this approach does not limit the practitioners in their choice of the reconstruction algorithm, and opens the possibility for researchers to also use the methods in research of novel reconstruction algorithms. But most importantly, simulating in projection space means the models need to process and produce an order-of-magnitude less data points, as compared to simulating the final reconstruction volumes. That means, simulating data points instead of . In our case, and . In practice, stays approximately the same, but is usually 3 – 4 larger.
@@ 3.1 Additive noise (baseline)
Before moving on to the development of more complicated methods, we wanted to see how a very simple method involving tilt-dependent scaling and the addition of Gaussian noise would approximate the target projections. The modality of projections produced by this method will be referred to as baseline projections in the rest of the text. First, we shifted and scaled each tilt of the noiseless input projections separately according to the average mean and standard deviation of each tilt within the training set. Through this tilt-dependent scaling, we attempt to model different degrees of attenuation as a function of tilt angle - the more extreme the tilt angle, the greater the attenuation due to the longer electron beam trajectory.
Next, we simply add Gaussian noise, as this is the natural choice of practitioners when they need to quickly create some simulated tilt-series. Of course, this raises the question of how much noise to add. In reality, practitioners would probably visually choose a value for that gives a similar signal-to-noise ratio. However, we opted for a more objective procedure by subtracting the noiseless content from the target projections to extract the noise and measure its statistics. In the case of SHREC data, where we know the ground truth, this task is easier than when using real data, as in this case we do not have the ground truth for the entire tomogram. So in a real scenario we would have to select regions for which we can estimate the ground truth and calculate the noise statistics only for these regions instead of using the whole tilt values. After adding the Gaussian noise, the resulting volume was scaled to match the average mean and standard deviation of the training set with respect to tilt. The steps for creating projections and consequently reconstructions using this method are shown in Figure 5, and an example of a simulated projection can be seen in Figure 3. The calculation of this modality takes only a few seconds with a single CPU.
@@ 3.2 Neural Style Transfer (faket)
To capture the noise structure of the benchmark projections more closely, in faket we implemented a more elaborate method of estimating the noise statistics as opposed to the simpler one used in baseline. In this case, we estimate the noise statistics for each tilt in the whole training set separately and fit a second-degree polynomial over the averages. This process captures the average noise statistics as a function of tilt angle based on the information from the whole training set. After adding the better estimated Gaussian noise, we obtained projections which we will refer to as noisy projections in this section. To further adapt the noisy projections we used the Neural Style Transfer technique introduced in Gatys et al. [2016].
NST was built to render the semantic content of natural images in different styles. At its core lies VGG net Simonyan and Zisserman [2014], a convolutional neural network optimized for object recognition and -localization. Within the NST framework, the VGG net is used for extracting the content and style representations of so-called “content” and “style” images, which are provided as inputs. NST then iteratively updates the output image to simultaneously match the content representation and the style representation of the provided inputs at multiple scales. NST is described in the aforementioned paper in great detail and we encourage the reader to consult it if any questions would arise.
@@ 4 Evaluation
To evaluate the quality of the proposed simulated projections and subsequently tomograms, no widely accepted metrics exist Seemann et al. [2020], and it is not clear how well the standard metrics such as the mean squared error, or more elaborate metrics such as the Fréchet inception distance Heusel et al. [2017], relate to the performance on tasks of interest to the practitioners.
@@
To obtain the main results, we trained DF for 70 epochs on benchmark, faket, and baseline training data, every time across 6 different seeds of randomness ( times in total). Each model was tested on the same test tomogram – model_9 from the original SHREC data set. The final comparison was done by observing the models’ performances in terms of F1 score for localization task and F1 macro score for classification task. The computation of the scores was done the same way as in Gubins et al. [2021] where also further particularities of the tasks are described. To estimate the 68% and 95% confidence intervals, we used the bootstrapping method. The results are visualized and discussed in Section 5.
@@ 4.1 DeepFinder
DeepFinder emerged from the SHREC challenge as one of the most successful candidates. One other reason why we chose to use DF was the availability of its source code. DF is a deep 3D-convolutional neural network trained in a supervised fashion by optimizing a dice loss. For training of the models, we used the Adam optimizer with learning rate of . The exponential decay rate was set to for the first moment estimate and for the second moment estimate as in Moebel et al. [2021]. To train the DF, we used tomograms containing particle instances from 16 classes (two of those being background and vesicles). We adjusted the training procedure of DF such that in one epoch, all patches containing particles within the training data set were seen once, or with minimal repetition. a patch in a shape of a box is rotated by 180∘ at random. Patches are processed in batches constructed from a randomly permuted list of all particles in the training set. After our changes in the training procedure, an epoch represents a different number of gradient steps than in Moebel et al. [2021], therefore the numbers of epochs are not comparable with the original paper. To perform the particle localization and -classification on the test tomogram, DF carries out a segmentation procedure followed by clustering and thresholding steps, where the two latter steps are computed using a CPU. The thresholding is used to reduce the number of false positive findings.
To train the DF implemented in Keras framework, we used multiple NVIDIA A100 40GB SXM4 GPUs. That allowed us to run multiple experiments in parallel. One training job comprising of 9 training tomograms required approx. 17GB of GPU memory, therefore we were able to submit two jobs on one GPU at a time. One training epoch, as defined earlier, took 21 min. to finish. The segmentation of a test tomogram took 2.5 min. on the aforementioned GPU. The clustering step took 30 min. on a single core of a 64-core AMD EPYC 7742 CPU. The most computational expensive step was thus the training of the DF.
@@ 5 Results
Comparison of DeepFinder’s performance on localization and classification tasks as a function of training data is presented in Table 1. The mean performances per epoch with 68% and 95% confidence intervals are presented in Figure 6 and Figure 7. Additional comparisons can be found in Appendix D.
@@
The performance of DF trained on benchmark data turned out to be the best in both localization and classification tasks, as expected, since the testing tomogram was reconstructed from the exactly same tilt-series as benchmark training set. In practice, however, it is infeasible to have such favorable conditions as fully labelled tomograms are not existent. Most biological studies focus on one or a subset of molecules to be labelled. And since the labelling process is laborious and requires domain experts, it is also expensive. With that in mind, we can hardly expect to have a training set of tomograms from exactly the same distribution as the tomograms we are interested in. Nevertheless, this result serves us, as the name suggests, as a benchmark that we are trying to reach with our proposed methods. The highest benchmark F1 score of 81.5% (68% CI [81.1%, 82.0%]) on localization task and 58.1% (68% CI [57.9%, 58.3%]) on classification task on average across 6 different random seeds was achieved after 654 epochs of training.
@@
Our fastest method of simulating the projections and subsequently the tomograms, baseline, scored 81.3% (68% CI [81.1%, 81.6%]) on localization task. It was a surprise that such a simple method led to so high localization performance, considering that it requires hours to create the benchmark projections, while it takes almost no time to create the baseline projections. This method can be therefore regarded as the “poor-man’s” choice in settings where the computational budget is very limited, or in settings where this task has to be performed repeatedly many times. Unfortunately, the classification performance of 44.1% (68% CI [43.2%, 45.1%]) was rather poor, not surprisingly, as classification is an inherently harder task, cf. Gubins et al. [2021].
@@
Our proposed method faket, based on additive noise (using our advanced estimation of in-projections noise) and subsequently neural style transfer, with on average 80.0% (68% CI [79.8%, 80.2%]) F1 score on localization task performed on par with the benchmark and significantly outperformed standard template matching algorithms. The best performance on the classification task was achieved after 654 epochs with a score of 53.3% (68% CI [52.8%, 53.8%]). With this result, faket matched the benchmark to 92 % while reducing the cost of data generation by a factor of 750 and using 33 times less memory. Multi-class classification performance certainly can not be reduced only to one number and it is important to consider also other metrics, such as per-class classification report and confusion matrices. The per-class classification performance is shown in Figure 8 and we provide the confusion matrices in Appendix A. The full performance reports are available in the accompanying repository for the most enthusiastic readers.
@@ 6 Conclusions
In this paper, we proposed faket, a method for simulating the forward operator of an electron microscope based on additive noise and Neural Style Transfer. The proposed method can be used for generating synthetic cryo-electron tilt-series which can be used to train deep neural networks to solve tasks such as particle localization and (much more challenging) particle classification. The field of cryo-electron tomography currently suffers from the lack of sufficient amounts of annotated data, and the proposed method aims to solve this problem. faket is a surrogate model that mimics the behavior of the physics-based TEM simulator SHREC, while being drastically less computationally expensive, both in terms of time and memory. It accelerates the data generation process by a factor of 750 while using 33 times less memory, making the generation of thousands of tilt-series feasible. Moreover, it is capable of simulating large tilt-series, which are common in experimental environments. For example, we generated a tilt-series on a single NVIDIA A100 40GB SXM4 GPU in less than 10 minutes. It also does not require any calibration protocol, it does not need to be retrained to be used on new data, nor does it require labeled reference data.
@@
The quality of our approximations was evaluated using the DeepFinder network, which emerged from the SHREC challenge as one of the most successful. The results showed that the performance of models trained using our approximations is on par with the benchmark method on localization task and reached 92 % of its performance on classification task while significantly outperforming standard template matching algorithms. This advance makes it possible to train particle localization and classification networks from scratch, or to pre-train networks that are later fine-tuned using manually labelled experimental data. The method is open source, and the experiments are reproducible.
@@
In future work, we will focus on further validation of this method using experimental data from TEM and we hope this advancement will serve as the basis for more software development in cryoET. Additionally, we plan to improve the method by using state-of-the-art neural networks and we want to provide the community of practitioners with a fully functional and easy-to-use piece of software to generate their synthetic data sets based on chosen particles from Protein Data Bank. The goal is to enable more accurate and efficient data analysis while also making the process more accessible to researchers in the field.
Appendix A Details on the performance for particle classification

### v2

@@ 3 Methods
The methods described in this section represent surrogate models that mimic the behaviour of the aforementioned SHREC physics-based simulator. The methods are to be applied in the projection space, i.e. all of them take noiseless (and ice-less) projections as inputs, shown in Figure 2, and produce real-like looking projections by matching their appearance with a target “style” projections. The outputs are ready to be reconstructed into final tomogram volumes using a reconstruction algorithm of choice. There are multiple reasons why we decided, and why it is reasonable, to model the simulator’s behaviour in projection space. Firstly, different reconstruction algorithms produce different types of artifacts and are suitable in different situations. Therefore, simulating only projections means the models do not have to mimic the artifacts crated by the reconstruction algorithm (e.g. smearing due to the missing wedge). Moreover, this approach does not limit the practitioners in their choice of the reconstruction algorithm, and opens the possibility for researchers to also use the methods in research of novel reconstruction algorithms. But most importantly, simulating in projection space means the models need to process and produce an order-of-magnitude less data points, as compared to simulating the final reconstruction volumes. That means, simulating data points instead of . In our case, and . In practice, stays approximately the same, but is usually 3 – 4 larger.
@@ 3.1 Additive noise (baseline)
Before moving on to the development of more complicated methods, we wanted to see how a very simple method involving tilt-dependent scaling and the addition of Gaussian noise would approximate the target projections. The modality of projections produced by this method is referred to as baseline projections in the whole text. First, we shifted and scaled each tilt of the noiseless input projections separately according to the average mean and standard deviation of each tilt within the training set. Through this tilt-dependent scaling, we attempt to model different degrees of attenuation as a function of tilt angle - the more extreme the tilt angle, the greater the attenuation due to the longer electron beam trajectory.
Next, we simply add Gaussian noise, as this is the natural choice of practitioners when they need to quickly create some simulated tilt-series. Of course, this raises the question of how much noise to add. In reality, practitioners would probably visually choose a value for that gives a similar signal-to-noise ratio. However, we opted for a more objective procedure by subtracting the noiseless content from the target projections to extract the noise and measure its statistics. With SHREC data, where precise ground truth is available, this task is less challenging than with real data, where we lack ground truth for the entire tomogram. So in a real scenario we would have to select regions for which we can estimate the ground truth and calculate the noise statistics only for these regions instead of using the whole tilt values. After adding the Gaussian noise, the resulting volume was scaled to match the average mean and standard deviation of the training set with respect to tilt. The steps for creating projections and consequently reconstructions using this method are shown in Figure 5, and an example of a simulated projection can be seen in Figure 3. The calculation of this modality takes only a few seconds using a single CPU.
@@ 3.2 Neural Style Transfer (faket)
To capture the noise structure of the benchmark projections more closely, in faket we devised a more elaborate method of estimating the noise statistics as opposed to the simpler one used in baseline. In this case, we estimate the noise statistics for each tilt in the whole training set separately and fit a second-degree polynomial over the averages. This process captures the average noise statistics as a function of tilt angle based on the information from the whole training set. After adding the better estimated Gaussian noise, we obtained projections which we will refer to as noisy projections. To further adapt those projections, we used the Neural Style Transfer technique implemented in PyTorch framework Paszke et al. (2019) and introduced in Gatys et al. (2016).
NST was built to render the semantic content of natural images in different styles. At its core lies VGG net Simonyan and Zisserman (2014), a convolutional neural network optimized for object recognition and -localization. Within the NST framework, the VGG net is used for extracting the content and style representations of so-called “content” and “style” images, which are provided as inputs. NST then iteratively updates the output image to simultaneously match the content representation and the style representation of the provided inputs at multiple scales. NST is described in the aforementioned paper in great detail and we encourage the reader to consult it if any questions would arise.
@@ 4 Evaluation
To evaluate the quality of the proposed simulated projections and subsequently tomograms, no widely accepted metrics exist Seemann et al. (2020), and it is not clear how well the standard metrics such as the mean squared error, or more elaborate metrics such as the Fréchet inception distance Heusel et al. (2017), relate to the performance on tasks of interest to the practitioners.
@@
To obtain the main results, we trained DF for 70 epochs on benchmark, faket, and baseline training data, every time across 6 different seeds of randomness ( times in total). Each model was tested on the same test tomogram – model_9 from the original SHREC data set. The final comparison was done by observing the models’ performances in terms of F1 score for localization task and F1 macro score for classification task. The computation of the scores was done the same way as in Gubins et al. (2021) where also further particularities of the tasks are described. To estimate the 68 % and 95 % confidence intervals, we used the bootstrapping method. The results are visualized and discussed in Section 5. Results of additional experiments are presented in the Appendix.
@@ 4.1 DeepFinder
DeepFinder emerged from the SHREC challenge as one of the most successful candidates. One other reason why we chose to use DF was the availability of its source code. DF is a deep 3D-convolutional neural network trained in a supervised fashion by optimizing a dice loss. For training of the models, we used the Adam optimizer with learning rate of . The exponential decay rate was set to for the first moment estimate and for the second moment estimate as in Moebel et al. (2021).
To train the DF, we used tomograms containing particle instances from 16 classes (two of those being background and vesicles). We adjusted the training procedure of DF such that in one epoch, all patches containing particles within the training data set were seen once, or with minimal repetition. a patch in a shape of a box is rotated by 180∘ at random. Patches are processed in batches constructed from a randomly permuted list of all particles in the training set. After our changes in the training procedure, an epoch represents a different number of gradient steps than in Moebel et al. (2021), therefore the numbers of epochs are not comparable with the original paper. To perform the particle localization and -classification on the test tomogram, DF carries out a segmentation procedure followed by clustering and thresholding steps, where the two latter steps are computed using a CPU. The thresholding is used to reduce the number of false positive findings.
To train the DF network implemented in Keras framework Chollet et al. (2015) using TensorFlow backend Abadi et al. (2015), we utilized multiple NVIDIA A100 40GB SXM4 GPUs. That allowed us to run multiple experiments in parallel. One training job comprising of 9 training tomograms required approx. 17GB of GPU memory, therefore we were able to submit two jobs on one GPU at a time. One training epoch, as defined earlier, took 21 min. to finish. The segmentation of a test tomogram took 2.5 min. on the aforementioned GPU. The clustering step took 30 min. on a single core of a 64-core AMD EPYC 7742 CPU. The most computational expensive step was thus the training of the DF.
@@ 5 Results
Comparison of DeepFinder’s performance on localization and classification tasks as a function of training data is presented in Table 1. The mean performances per epoch with 68 % and 95 % confidence intervals (CI) are presented in Figure 6 and Figure 7. Results of additional experiments can be found in Appendices C, LABEL:, D, LABEL: and E.
@@
The performance of DF trained on benchmark data turned out to be the best in both localization and classification tasks, as expected, since the testing tomogram was reconstructed from the exactly same tilt-series as benchmark training set. In practice, however, it is infeasible to have such favorable conditions as fully labelled tomograms are not existent. Most biological studies focus on one or a subset of molecules to be labelled. And since the labelling process is laborious and requires domain experts, it is also expensive. With that in mind, we can hardly expect to have a training set of tomograms from exactly the same distribution as the tomograms we are interested in. Nevertheless, this result serves us, as the name suggests, as a benchmark that we are trying to reach with our proposed methods.
The highest benchmark F1 score of 81.5 % (68 % CI [81.1 %, 82.0 %]) on localization task and 58.1 % (68 % CI [57.9 %, 58.3 %]) on classification task on average across 6 different random seeds was achieved after 654 epochs of training.
To put these numbers in perspective, we investigated the limits of DF performance by training and evaluating on completely noiseless data, see Appendix D for more details. In this setting DF achieved F1 score of 83.2 % (68 % CI [82.4 %, 84.0 %]) on localization task and 72.5 % (68 % CI [70.5 %, 74.6 %]) on classification task.
@@
Our fastest method of simulating the projections and subsequently the tomograms, baseline, scored 81.3 % (68 % CI [81.1 %, 81.6 %]) on localization task. It was a surprise that such a simple method led to so high localization performance, considering that it requires to create the benchmark projections, while it takes almost no time to create the baseline projections. This method can be therefore regarded as the “poor-man’s” choice in settings where the computational budget is very limited, or in settings where this task has to be performed repeatedly many times. Unfortunately, the classification performance of 44.1 % (68 % CI [43.2 %, 45.1 %]) was rather poor, not surprisingly, as classification is an inherently harder task, cf. Gubins et al. (2021).
@@
Our proposed method faket, based on additive noise (using our advanced estimation of in-projections noise) and subsequently neural style transfer, with on average 80.0 % (68 % CI [79.8 %, 80.2 %]) F1 score on localization task performed on par with the benchmark and significantly outperformed standard template matching algorithms. The best performance on the classification task was achieved after 654 epochs with a score of 53.3 % (68 % CI [52.8 %, 53.8 %]). With this result, faket matched the benchmark to 92 % while reducing the cost of data generation by a factor of 750 and using 33 times less memory. Multi-class classification performance certainly can not be reduced only to one number and it is important to consider also other metrics, such as per-class classification report and confusion matrices. The per-class classification performance is shown in Figure 8 and we provide the confusion matrices in Appendix E. The full performance reports are available in the accompanying repository for enthusiastic readers.
In Appendix C, we present further experimental results showcasing the fine-tuning of faket models using a subset of benchmark data. This approach further narrows the performance gap between the faket and benchmark models, achieving 97 % of the benchmark model’s classification performance, while exceeding its performance on localization task. These results might be valuable for cryoET practitioners who are seeking to maximize the performance of their models and have, or can get a small but representative sample of their data labeled.
@@ 6 Conclusions
In this paper, we proposed faket, a method for simulating the forward operator of an electron microscope based on additive noise and Neural Style Transfer. The proposed method can be used for generating synthetic cryo-electron tilt-series which can be used to train deep neural networks to solve tasks such as particle localization and (much more challenging) particle classification. The field of cryo-electron tomography currently suffers from the lack of sufficient amounts of annotated data, and the proposed method aims to solve this problem. faket is a surrogate model that mimics the behavior of the physics-based TEM simulator SHREC, while being drastically less computationally expensive, both in terms of time and memory. It accelerates the data generation process by a factor of 750 while using 33 times less memory, making the generation of thousands of tilt-series feasible. Moreover, it is capable of simulating large tilt-series, which are common in experimental environments. For example, we generated a tilt-series on a single NVIDIA A100 40GB SXM4 GPU in less than 10 minutes. It also does not require any calibration protocol, it does not need to be retrained to be used on new data, nor does it require labeled reference data. The method is open source, and the experiments are reproducible.
@@
The quality of our approximations was evaluated using the DeepFinder network, which emerged from the SHREC challenge as one of the most successful. The results showed that the performance of models trained using our approximations is on par with the benchmark method on localization task and reached 92 % of its performance on classification task while significantly outperforming standard template matching algorithms. When further fine-tuned using a portion of benchmark data, the classification performance was improved to 97 %. This advancement simplifies the generation of fully-labeled, high-quality synthetic tilt-series that resemble experimental TEM data requiring analysis. This simulated data can either be used to train particle localization and classification neural networks from scratch, or serve as fpre-training data for networks that will be fine-tuned with manually labeled experimental data later. Researchers investigating reconstruction algorithms can also benefit from our simulator since the availability of ground truth allows for effortless evaluation of their novel methods.
@@
In future work, we will focus on further validation of our method using experimental data from TEM and we hope this advancement will serve as the basis for more method development in cryoET. Additionally, we plan to improve the method by using state-of-the-art neural networks and we want to provide the community of practitioners with a fully functional and easy-to-use piece of software to generate their synthetic data sets based on chosen particles from Protein Data Bank. The goal is to enable more accurate and efficient data analysis while also making the process more accessible to researchers in the field.
Appendix A Side-by-side comparison of projection types
In pursuit of re-creating the benchmark reconstructions to match those from SHREC data set as closely as possible, we inspected the public source code of PyTom Hrabe et al. (2012) for clues on which filters might have been used. In the original paper, authors only mention that they use weighted back-projection. Which would suggest the use of a simple ramp filter. In a personal communication, they mentioned using a default filter from their private PyTom version with undefined Crowther frequency. Therefore, we implemented and used a simple ramp filter with a support for Crowther frequency. Then, we visualized SHREC’s and our reconstructions in Fourier space to visually compare the results. Unfortunately, none of the variations of a simple ramp filter, even with Crowther frequency, produced the expected result similar to SHREC. For that reason, we started to iteratively change our filter in order to approximately match the aforementioned reconstructions in both real and Fourier space, in hope of finding the desired match. After many iterations, we found a good-enough match, a filter which is a product of a 2D Gaussian filter (, ), a 1D ramp filter broadcasted to 2D ( Nyquist), and a 2D circular filter (, i.e. Nyquist frequency). Exact implementation is available in code repository.
Appendix C Further improving the performance of DF models trained with faket data by subsequent fine-tuning on a fraction of benchmark data
Fine-tuning is a prevalent approach in machine learning where a pre-trained model is further refined using a subset of the available data specific for the task. This methodology often yields improved performance by allowing the model to specialize on the nuances of a specific dataset. For practitioners in cryoET, this process can be of great value when seeking to maximize the performance of their models. If they are able to label a small but representative sample of their data manually, they can leverage this technique to their advantage.
Here, we compared the performance of models trained on baseline or faket data with that of a model initially trained on faket data and later fine-tuned using a subset (3 out of 9) of benchmark tomograms, denoted as finetuned-3 in the figures. The fine-tuning began after the 50th epoch, triggering a notable improvement in both localization and classification performance. To make a reasonable comparison, we include results from a model, benchmark-3, trained solely on the same 3 benchmark tomograms used for fine-tuning.
During the fine-tuning and benchmark-3 training stages, the total number of training tomograms was smaller. To maintain experiment consistency by keeping an approximately equivalent number of gradient steps per epoch, we adapted our definition of an epoch. Here, it refers to a cycle where the model is trained on 9 tomograms — meaning that during fine-tuning, three full training cycles with 3 tomograms each are considered one epoch.
Appendix D Additional experiments
For our baseline, we chose the conventional approach of adding Gaussian noise. However, acknowledging the noise component in TEM projections fluctuates with tilt angle, we devised a tilt-dependent algorithm for Gaussian noise addition. This formed the basis for our NST-based faket method, producing what we refer to as noisy projections. One might wonder if such simulation enhances DF performance, a query best illustrated by a figure. Although localization performance remained consistent, we observed a slight enhancement in classification, as shown in Figure 13 (top). Despite this modest improvement, we didn’t deem it significant enough to modify our simple baseline method, but still opted to incorporate it into our faket method.
We present an ablation study featuring a faket-random model trained on data created with a randomly initialized NST network, in contrast to pre-trained. Results in Figure 13 (middle) underscore the necessity of pre-training for successful simulation and suggest that pre-training using cryoET data might further enhance the simulated data’s resemblance to real tomograms — a question still under exploration.
Finally, to offer insight into the potential limits of DF on the studied data, we share results from a noiseless experiment in Figure 13 (bottom). The model was trained and tested on completely noiseless reconstructions, making it also the only model in this paper evaluated on a noiseless test tomogram.
Appendix E Further details on particle classification performance